Im confused on how to do rational equations and sadly its urgent this paper is due tonight so please help im desperate!!!!!!! My problem is: n^2/2 - 5n/6 - 2 = 0
Answer by Alan3354(21605)

n^2/2 - 5n/6 - 2 = 0
3n^2 - 5n - 12 = 0
(3n + 4)*(n -3) = 0
n = 3
n = -4/3